a rectangle is 7 cm longer than it is wide. its perimeter is at least 58 cm. what are the smallest possible dimensions for the rectangle
Answer by partha_ban(41)

Let w = width
Then length = w + 7
Therefore, 2(length + width) = 58
2(w+7+w) = 58
w + 7 + w = 58/2 = 29
2w + 7 = 29
2w = 29 - 7 = 22
w = 22/2 = 11
Therefore width = 11 cm and length = 11+7 = 18 cm